OpenStudy (anonymous):

PLEEEEEAAAAAAAAASSSSE HELP? :( (re-post) A job pays 32000$ for the first year with an annual increase of 6% per year beginning in the second year. What is the salary in the sixth year? What is the total salary paid over this six-year period?

OpenStudy (anonymous):

take 32,000*1.06=33,920 will give u the salary in the second year, do the same thing ofr third year by take the salary in second years times 1.06, so 33,920*1.06=35,955.2 (third year) 35955.2*1.06= 38,112.5 will give u the fourth year salary.

OpenStudy (anonymous):

do the same thing until 6th year.

OpenStudy (anonymous):

i need a formula

OpenStudy (anonymous):

like using a sum formula

OpenStudy (anonymous):

i was thinking 32,000(0.06*x+1), but it doesn't seem right.

OpenStudy (anonymous):

i have one that says sn=a1(1-r^n)/1-r but i always get the wrong answer :/

OpenStudy (anonymous):

that's can't be right

OpenStudy (anonymous):

where did u get that formula from?

OpenStudy (anonymous):

my teacher - i could have written it wrong - wait

OpenStudy (anonymous):

32000 (1 - 0.06^n)/(1 - 0.06) this is right.

OpenStudy (anonymous):

thank you, another question why do i keep getting 31999.93?

OpenStudy (anonymous):

rounding error.

OpenStudy (anonymous):

shouldn't the real answer be 4282322?

OpenStudy (anonymous):

nope, does it make sense to have that high of salary? how many years did u put in there?

OpenStudy (anonymous):

6 because it asked for the sixth year

OpenStudy (anonymous):

nope, 6 should be n=5, because the first year is 0, and second year is 1

OpenStudy (anonymous):

u think you would make that much of money in 6 years? it doesn't make any sense to get that high salary after 6 years

OpenStudy (anonymous):

ohhhhhhhhhh, ok thanks i didn't notice that

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

no problemo.